OpenStudy (16schro_alan):

An unopened cereal box contains 15/28 cubic foot of cereal. If the box is full, and the length and height of the box are 5/7 and 1 1/2 feet, respectively, what is the width of the box?

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

volume is the product of all three dimensions you have to solve \[\frac{5}{7}\times 1\tfrac{1}{2}\times x=\frac{15}{28}\] for \(x\)

OpenStudy (misty1212):

first off, do you know how to compute \[\frac{5}{7}\times 1\tfrac{1}{2}\]?

OpenStudy (16schro_alan):

no im not sure

OpenStudy (misty1212):

turn \(1\tfrac{1}{2}\) in to an improper fraction, then multiply straight across

OpenStudy (misty1212):

so i guess the next question is, do you know what \[1\tfrac{1}{2}\] looks like as a fraction?

OpenStudy (16schro_alan):

idk 1.5?

OpenStudy (misty1212):

lol that is not a fraction, that is a decimal

OpenStudy (misty1212):

a fraction has a numerator and a denominator

OpenStudy (misty1212):

i can show you if you like

OpenStudy (16schro_alan):

3 1/2

OpenStudy (misty1212):

no, one and a half is not three and a half

OpenStudy (misty1212):

\[1\tfrac{1}{2}=\frac{2\times 1+1}{2}\]

OpenStudy (amour.anohs):

She Ment Multiply The Whole Number By The Denominator And Add What You Got Plus The Numerator .After You Do That Use What You Got When You Added And Multiplied As The Numerator ,And Leave The Denominator The Same. @misty1212

OpenStudy (amour.anohs):